OpenStudy (anonymous):

medal!! what is the median for the set of numbers? 1, 3, 2, 1, 1, 2, 1, 4, 1, 2, 2, and 1.

OpenStudy (anonymous):

It's 2. That's what is in the middle of the data set when you organize them in numerical order.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

acutally it is the average isn't it 1.83

OpenStudy (anonymous):

but wouldn't there be 2 numbers in the middle

OpenStudy (campbell_st):

well can you arrange the data in order... smallest to largest

OpenStudy (anonymous):

I thought that the median was the number in the middle, and theres 2 middle numbers so what do I do?

OpenStudy (campbell_st):

then there are 12 pieces of data so the position of the data is (12 +1)/2 = 6.5 the median is between the 6th and 7th score

OpenStudy (anonymous):

what?......

OpenStudy (campbell_st):

then the median is Median = (6th score + 7th Score)/2

OpenStudy (anonymous):

ooooohhh so since there are 2 middle numbers then I would have to do a number in between them like 6.5?

OpenStudy (campbell_st):

well after arranging the data in order.... the easy way to find the position of the median is 1. count how many scores 12 2. add 1 to the number 12 + 1= 13 3. divide the answer by 2 13/2 = 6.5 which tells you the position of the median is between the 6th and 7th scores...

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

WOW!!! THANK U!!!

OpenStudy (campbell_st):

so the 6th score is 1 and the 7th score is 2 so the median is median = (1 + 2)/2

OpenStudy (anonymous):

so 1.5